The third angle of an isosceles triangle is 8° more than twice as large as each base angle. Find the measure in degrees of each base angle?​
Answer by ikleyn(52776)

.
The third angle of an isosceles triangle is 8° more than twice as large as each base angle.
Find the measure in degrees of each base angle?​
~~~~~~~~~~~~~~~~~~~


Let x be the measure of a base angle.

Then the measure of the other base angle is x, again,
and the measure of the third angle (the "vertex" angle) is 2x+8 degrees.


Since the sum of interior angles of any triangle is 180 degrees, we have this equation


    x + x + (2x+8) = 180,


which implies


    4x + 8 = 180,

       4x  = 180 - 8 = 172

        x            = 172/4 = 43  degrees.


ANSWER.  The measure of each base angle is 43 degrees.

Solved.